OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (anonymous):

0=x,2=y 2=4(0)+b

OpenStudy (anonymous):

wait... perp, not parallel

OpenStudy (anonymous):

2=-1/4(0)+b

OpenStudy (anonymous):

y = –4x + 3 y = 4x + 2 y = -1/4x + 2 y = 1/4x + 2

OpenStudy (anonymous):

2=0+b 2=b y=-1/4 x+2

OpenStudy (anonymous):

wait-ur original line was 4x+3, right?

OpenStudy (anonymous):

The firts thing that you need to do is sketch the graph then second thing u must know is that for the lines to be parrallel their slopes hava to multiplied "-1"= m1*m2=-1 then you have to find a line function with slope of -1/4 that crosses the point(0,2)

OpenStudy (amistre64):

the first thing you need to do ..... realize that they give you slope intercept form, and give you a new intercept number and tell you to modify the slope to perp

OpenStudy (amistre64):

th3 perp of 4 is NOT 1/4 , ill assume thats a typo tho

OpenStudy (amistre64):

just like mines a typo lol

OpenStudy (anonymous):

what??

OpenStudy (amistre64):

your slope of 1/4 is in error. the slope of a perpendicular lines have a product of -1 4 * (1/4) NOT= -1

OpenStudy (amistre64):

4*p = -1 p = -1/4 is the slope of any perp line to the original equation

OpenStudy (anonymous):

u r absolutely right.i have done a mistake

OpenStudy (anonymous):

y= -1/4x+2